# Dynamic Adversarial Attacks on Autonomous Driving Systems

### v1

@@ 5.2 Attack Procedure
Step 1. We collect images from the display screen to construct the SIT-Net as explained in Section 4.2.
@@
Step 2. We run robot cars in each scenario to collect synchronized Lidar and image data. This dataset is clustered as explained in Section 4.1. The number of clusters hinges on the total distance traversed and the precision of positional estimations. For our experiments, we set the cluster count to three. We utilize this processed dataset to optimize a set of patches using Algorithm 1. We set the training epoch at 1000 for each cluster and start with a randomly generated initial patch. Concurrently, a static patch is trained across all clusters for an identical epoch count.
@@
Step 3. After training, we evaluate the effectiveness of the generated patches by the attack success rate on a subset including images captured in a similar environment and background and on a dataset of images with a background that is not in the attack dataset. An attack is considered successful if the patch makes the object detector predict a stop sign with confidence greater than 0.5 in place of the original sign.
5.3 Results and Discussion
@@
The loss over iterations during training of the patch for each cluster and all dataset is depicted in Figure 7. This demonstrates that optimizing adversarial patches specifically for individual data clusters significantly expedites the training process and enhances the overall effectiveness of the attack. This cluster-based approach benefits from the inherent homogeneity within each cluster, allowing for more rapid optimization compared to a normal training approach on a broader dataset. Dynamic patches, tailored to specific environmental conditions and scenarios, clearly outperformed static patches that attempted to cover all data variances. This is visually represented in Figure 8, which illustrates the variations in attack strategies. A video showing attacks using dynamic patches in various autonomous driving scenarios involving two R2 robot cars is available online [^2] https://youtu.be/2bLEHNTJ1pI. In addition, Figure 9 shows the optimized patches for different clusters and the entire dataset.
@@
Moreover, the designed objective function demonstrates the efficiency in generating a stop sign in the location of the original traffic sign as shown in Figure 8. This functionality is essential for the success of the adversarial attack, as it directly influences the decision-making algorithms of the autonomous driving systems being attacked. By misclassifying crucial traffic signs, the dynamic patches effectively mislead the autonomous vehicle, compelling it to make erroneous navigational decisions.
Further, Table 1 reveals that the success rate of the attacks is higher with the dynamic patch in most cases. This is because the chance of deception is higher in closer clusters. This is significant because it aligns well with the practical requirements of autonomous driving systems, where the accuracy and reliability of object detection are most critical at closer distances for immediate navigational decisions. The enhanced success rate in closer clusters means that our approach could potentially be more effective in real-world scenarios, where misclassification of traffic signs at closer ranges poses a higher risk. Consequently, this could lead to more effective strategies for reducing risks associated with the navigation and planning of autonomous vehicles.
5.4 Conclusions
@@
We formulate a novel approach for attacking object detectors by employing dynamic patches displayed on a screen to influence the decisions of autonomous driving systems. We have demonstrated that data clustering substantially accelerates the training process and enhances the overall efficacy of the attack. Additionally, to address the issue of color transformation in captured images of the displayed patch, we have introduced a straightforward CNN model capable of compensating for environmental factors. Our findings underscore the potential vulnerabilities in the perception mechanisms of autonomous driving systems and highlight the importance of considering such adversarial tactics in the development of robust decision-making algorithms.

### v2

@@ V-B Attack Procedure
Step 1. We collect images from the display screen to construct the SIT-Net as explained in Section IV-B.
@@
Step 2. We run robot cars in each scenario to collect synchronized Lidar and image data. This dataset is clustered as explained in Section IV-A. The number of clusters hinges on the total distance traversed and the precision of positional estimations. For our experiments, we set the cluster count to three. We utilize this processed dataset to optimize a set of patches using Algorithm 1. We set the training epoch at 1000 for each cluster and start with a randomly generated initial patch. Concurrently, a static patch is trained across all clusters for an identical epoch count.
@@
Step 3. After training, we evaluate the effectiveness of the generated patches by the attack success rate in an extensive number of experiments including a similar environment and background and in settings with a background that is not in the attack dataset. For each of the sign types, approximately 1000 image frames are evaluated. An attack is considered successful if the patch makes the object detector predict a stop sign with confidence greater than 0.5 in place of the original sign.
We expand our evaluation to include comparisons with methodologies similar to ours, particularly focusing on sticker-based attacks as explained in [34], which represent the closest parallel in existing literature. Unlike previous studies where patches are directly attached to the target object, in our approach, the adversarial patch and the target object are positioned separately. This unique setup necessitated the optimization of our patches through Equation III-C coupled with the consideration of the non-printability score, which is elaborated in [34, 28].
Moreover, we conduct experiments in which no patch was displayed to have a benign case. Table I presents the results of these experiments.
Attack Method Dynamic (%) Static (%) Printed (%) Benign (%)
Go-straight sign 39.9 30.0 1.3 0.0
Turn sign 27.4 22.1 3.0 0.0
Pedestrian sign 18.0 15.2 2.0 0.0
TABLE I: Comparison of attack success rates for Dynamic, Static, and Printed Patches across different sign types in an intersection scenario. This table also shows the control group (“Benign”) where no adversarial attack was present, indicating no false detections under benign conditions.
Figure 8: Display of trained adversarial patches across various clusters and the entire dataset. The rows represent different traffic sign scenarios, specifically “Go-straight”, “Turn”, and “Pedestrian” signs, in that order. For the columns, from left to right, we showcase patches optimized for clusters based on proximity to the target: “Near”, “Mid”, and “Far”, followed by patches trained on the combined dataset (denoted as ‘All’). This arrangement illustrates how the patches are tailored to specific distances from the target object.
(a) Confidence distribution for misclassifying ”Go-straight” sign as ”Stop-sign”
(b) Confidence distribution for misclassifying ”Turn” sign as ”Stop-sign”
(c) Confidence distribution for misclassifying ”Pedestrian” sign as ”Stop-sign”
(d) Confidence distribution for misclassifying ”Go-straight” sign as ”Stop-sign” using a printed patch
(e) Confidence distribution for misclassifying ”Turn” sign as ”Stop-sign” using a printed patch
(f) Confidence distribution for misclassifying ”Pedestrian” sign as ”Stop-sign” using a printed patch
Figure 9: Comparative histograms illustrating the confidence levels in misclassifying ”Go-straight”, ”Turn”, and ”Pedestrian” signs as ”Stop-sign” under two conditions: with dynamic/static patches (top row) and with printed adversarial patches (bottom row). These graphs highlight the effectiveness and variation of the adversarial attacks across different scenarios, noting that higher confidence values indicate better performance in the attacks. The spread and peaks of the confidence levels evidence the effectiveness of dynamic patches in deceiving.
(a) Effectiveness range of patches for the ”Go-straight” sign attack
(b) Effectiveness range of patches for the ”Turn” sign attack
(c) Effectiveness range of patches for the ”Pedestrian” sign attack
Figure 10: Scatter plots contrasting the effectiveness of dynamic and static patches in “Stop-sign” adversarial attacks as a function of the distance (in meters) of camera car to both the sign and the patch car. The dynamic patches (displayed in red-orange) demonstrate higher effectiveness across a more extensive range of distances from the camera to the patch, as indicated by the spread of the data points. The color intensity reflects the classifier’s confidence in misclassification, with the dynamic patch showing high confidence over larger distances compared to the static patch (shown in blue).
V-C Results and Discussion
@@
The loss over iterations during training of the patch for each cluster and all datasets is depicted in Figure 7. Also, Figure 8 shows the optimized patches for different clusters and the entire dataset. Figure 7 demonstrates that optimizing adversarial patches specifically for individual data clusters significantly expedites the training process and enhances the overall effectiveness of the attack. This cluster-based approach benefits from the inherent homogeneity within each cluster, allowing for more rapid optimization compared to a normal training approach on a broader dataset. The success rate in Table Iindicate that Dynamic patches consistently outperform Static and Printed patches that attempted to cover all clusters in the dataset. Printed patches show significantly lower effectiveness across all sign types. We discus the ineffectiveness of these printed patches in Appendix A-C. The benign case, with a 0.0% success rate, confirms the absence of false positives in normal conditions. Further, Figure 9 displays the confidence distribution for the misclassification of “Stop-sign” for three sets of adversarial patch attacks on the “Pedestrian”, “Turn”, and “Go-straight”. The dynamic patch consistently outperforms the static patch in every case in terms of confidence level. A visual depiction of the attack efficiency corresponding to the camera car’s distance from the patch car and the sign is shown in Figure10. In contrast to the static patch, which has a more concentrated cluster of effectiveness at shorter ranges, the dynamic patch’s effectiveness does not decrease as significantly with greater distance. Overall, Figure 9 and Figure 10 point to the fact that dynamic patches are more successful than static patches in deceiving the object detector over a variety of instances. This can be explained by the adaptive nature of dynamic patches, which are designed to take into account different perspectives and distances, making them more resilient to environmental changes. This is visually represented in Figure 12, which depicts the consecutive frames of the attacks. A video showing attacks using dynamic patches in various autonomous driving scenarios involving two R2 robot cars is attached. 222 https://youtu.be/Wh2sPYpWczQ
@@
Moreover, the designed objective function demonstrates the efficiency in generating a stop sign in the location of the original traffic sign as shown in Figure 12. This functionality is essential for the success of the adversarial attack, as it directly influences the decision-making algorithms of the autonomous driving systems being attacked. By incorrectly identifying key traffic signs, the dynamic patches deceive the autonomous vehicle into stopping or passing prematurely, leading to incorrect navigational choices.
V-D Ablation Study on the Effectiveness of SIT-Net
To evaluate the necessity and impact of the Screen Image Transformation Network (SIT-Net) within our adversarial framework, we conducted an ablation study where both static and dynamic adversarial patches were trained without incorporating SIT-Net. This experimental setup allowed us to isolate and study the specific contributions of SIT-Net, which simulates realistic environmental influences like color transformation and contrast adjustments that are crucial in real-world scenarios.
In our simulations, we overlaid the trained patches on images to mimic their real-world deployment on screens. The patches performed effectively under these ideal conditions, demonstrating their potential efficacy. However, when these patches were actually displayed on screens in a real-world environment, their effectiveness drastically declined, with a success rate dropping to zero. This significant discrepancy underscores the critical role of SIT-Net in preparing adversarial patches to cope with real-world variability and environmental conditions.
Furthermore, we found that simulations lacking SIT-Net do not adequately capture the visual characteristics of adversarial patches as displayed in real-world conditions. In simulations, the patches may appear sharper and have clearer colors than they do in real-world environments. This difference results in an overestimation of patch effectiveness during optimization, as the simulated visual quality exceeds what is typically achievable in practical deployments. This emphasizes the essential role of SIT-Net in providing a realistic preparation and evaluation of adversarial patches, ensuring that their performance under real-world conditions is accurately represented.
Figure 11 showcases the dramatic changes in patches when displayed on a screen, highlighting the importance of including SIT-Net in the training process to achieve realistic and effective adversarial patches in practical deployments.
Figure 11: Illustration of adversarial patch effectiveness without SIT-Net. The top row depicts real-world implementation results, and the bottom row shows simulation results. The columns, from left to right, represent attacks on “Go-straight”, “Turn”, and “Pedestrian” signs. This visual comparison clearly shows that the simulation without SIT-Net fails to replicate the real-world effectiveness of the patches, emphasizing the importance of SIT-Net in realistic adversarial training.
V-E Conclusions
@@
We formulate a novel approach for attacking object detectors by employing dynamic patches displayed on a screen to influence the decisions of autonomous driving systems. We have demonstrated that data clustering substantially accelerates the training process and enhances the overall efficacy of the dynamic attack. Additionally, to address the issue of color transformation in captured images of the displayed patch, we have introduced a straightforward CNN model capable of compensating for environmental factors.
Our findings underscore the potential vulnerabilities in the perception mechanisms of autonomous driving systems and highlight the importance of considering such adversarial tactics in the development of robust decision-making algorithms. Future work will focus on developing strategies to enhance the robustness of these systems against adversarial attacks. This includes exploring advanced machine learning techniques to detect and mitigate the effects of dynamic adversarial patches in real-time, as well as enhancing the training methods to incorporate a wider range of environmental variations.
(a) Adversarial patch attacks on the Go-straight sign.
(b) Adversarial patch attacks on the Turn sign.
Figure 12: Sequential frames depicting two types of adversarial patch attacks on an autonomous vehicle’s perception system. For each sub-figure, the top row shows the top view of a dynamic patch attack scenario. The middle row illustrates a dynamic patch attack, wherein the adversarial patches displayed on the screen change based on the distances to the camera and the sign, leading to a more deceptive misclassification across consecutive frames. The bottom row presents a static patch attack, where a single, unchanging patch is used. Both attacks aim to disrupt the perception system, causing it to mistakenly identify the signs as Stop signs.
Appendix A Appendix
A-A Dataset Overview
The dataset utilized for training our adversarial patches was meticulously compiled to ensure a broad representation of various scenarios encountered in autonomous driving. This dataset includes synchronized LiDAR and image data collected from our robot cars in indoor environments for each sign, which is critical for training both static and dynamic adversarial patches effectively. Figure 13 displays examples from our dataset, showing the diversity in the images. The full dataset is provided in this link [^3] https://drive.google.com/drive/folders/1UiODhj44Wos0TJAiK1067lCwvnoJt0qu?usp=sharing.
(a) Go-straight sign
(b) Turn sign
(c) Pedestrian sign
Figure 13: Example images from dataset for each sign, demonstrating the variety of scenarios in the dataset.
The dataset for each sign is divided into three distinct clusters based on the distance to the screen and sign as discussed in section IV-A2, which are pivotal for the effectiveness of dynamic patches in varying scenarios. These clusters are specifically designed to encapsulate different ranges of distances between the camera car and the target objects, ensuring that each patch is optimized for its respective range. The clustering approach helps in tailoring the patches to specific environmental conditions and object proximities. Figure 14 provides a detailed breakdown of the number of images in each cluster for each type of sign.
(a) Go-straight sign
(b) Turn sign
(c) Pedestrian sign
Figure 14: Distribution of images across the three clusters, showing the balance of clusters for each sign.
A-B Natural Lightening Condition Test
We conducted additional experiments to assess the robustness of our adversarial patches under various outdoor lighting conditions using patches that were initially trained in an indoor environment. Even though our adversarial patches were trained for indoor controlled conditions, they remained effective in outdoor lighting conditions. Figure 15 illustrates the performance of adversarial patches under different natural lighting conditions at various times of the day. These results highlight the resilience of our approach across a broad spectrum of natural lighting situations, though they also highlight areas for potential enhancement, particularly in very harsh lighting conditions.
Figure 15: Adversarial patch attacks in natural lighting conditions at different times of the day.
Additionally, to quantify the effectiveness more precisely, we generated scatter plots contrasting the performance of dynamic and static patches. Figure 16 provides a detailed view of the “Stop-sign” adversarial attacks and their effectiveness as a function of the distance (in meters) between the camera car and both the sign and the patch car in outdoor lighting conditions.
Figure 16: Scatter plots contrasting the effectiveness of dynamic and static patches in “Stop-sign” adversarial attacks as a function of the distance (in meters) of camera car to both the sign and the patch car in the outdoor lighting condition.
A-C Ineffectiveness of Printed Patches in Real-World Scenarios
We present the extended analysis of printed adversarial patches, highlighting the practical implications of the non-printability loss on real-world efficacy. The non-printability score, as discussed in [34], ensures that the colors utilized in the patches can be accurately reproduced by standard printers. However, this requirement often limits the overall effectiveness of the attack by restricting the range of colors that can be used. This constraint leads to a significant reduction in the effectiveness of the printed patches.
Figure 17 illustrates that the confidence levels for the “stop-sign” induced by these patches consistently fall below the critical threshold of 0.5. This outcome directly reflects the limitations imposed by the non-printability loss, which, while intended to make the patches printable, does not adequately mirror the appearance of real-world adversarial patches.
Figure 17: Sequential frames showing the performance of adversarial printed patch attacks. The frames depict that the confidence levels of the misclassified “Stop-sign” remain below the critical threshold of 0.5, indicating the ineffectiveness of these patches in real-world conditions.